You are here: Foswiki>Commissioning Web>Com20100629 (02 Jul 2010, PascaleHibon)Edit Attach

Commissioning: 29 June 2010 UT

Observer: Pascale H, Kate R, Dave T (LBTO)
Telescope Operator: David G (LBTO)
Telescope Support: (LBTO)
Instrument Support: Tim S (LBTO)
Adaptive Secondary Operator: Kate R (LBTO) Marco X (Firenze)

Summary

Lost time due to M2 astigmatism problem: 2 hrs

- Transformation data: StoneLagw data stopped after only 7 cubes due to OSS stopped due to trying to communicate with left side instead of right.
- Took preset data for elevations 85/70/55/30: Collimation was taking a very long time ~ 5 or more minutes due to excessive amounts of astigmatism. - AO team was called and decided to disengage the automatic astigmatism on mirror, this made collimation much better. - Took Preset data on BS9166 at EL= 45 and EL=32 - Took Guiding/ Collimation data on BS9177 EL =63, BS9169 EL=42, BS9188 EL=70. Aborted BS9188 due to loss of guide star. Tried on several other stars, unsuccessful due to bright sky and WFSC limits.

Preparations

Opened at 02:51 UT
AO shell required 4 stops and restarts of adsc system before wind speed, elevation, and dew point would properly update.

Details

03:13 pointto Arcturus TelMode="TRACK" gs=0 PA=0
03:16 Arcturus aquired after ptspiral
03:21 pointto BS9158 TelMode="ACTIVE" gs=0
03:26 pointto BS9158 TelMode="ACQUIRE" gs=0
03:28 pointto BS9158 TelMode="ACTIVE" gs=0: collimated, seeing ~0.67"
03:41 cd /OldRepository/IRTC_Data/20100627/transform
03:42 prepdir
03:42 pointto StoneLagw TelMode="ACTIVE" gs=0
03:47 collimated, seeing ~ 0.64 - 0.72"
03:54 pointto StoneLagw TelMode=ACTIVE gs=35PA=0
03:56 takepic 2.0
  • image: irtc.20100629.035623.fits
  • image in ds9 is too faint, retry with longer exposure.
03:58 takepic 3.0
  • image: irtc.20100629.035843.fits
04:04 offset 30 0
04:05 takeseq 3.0 20
  • image: irtc.20100629.040451 cube 65
04:07 offset 0 0
04:09 cl < tran_StoneL_mix1.cl
Seeing ~ 0.78
041034 cube 66
041241 cube 67
041448 cube 68
041653 cube 69
041857 cube 70
042136 cube 71
042342 cube 72

04:31 Script died (got stuck) due to OSS: David calls Paul to take a look. Script stuck on Offset: StoneLagw -> GS20
04:34 Pascale updated script from last position, created: tran_StoneL_mix1_mod.cl

04:42 Paul calls with error in OSS as the OSS was trying to communicate with left side instead of right when presets were sent.
04:54 Seeing ~ 1.0"-1.3"; abandoning the transformation data for now

Preset Data:

Seeing ~0.9"-1.2"
Pointing to stars at different Elevations: 85/70/55/40/30 etc... with BS stars; allowing it to collimate, then takepic 2.0 and recording image in log.
05:25 pointto BS9166 TelMode="ACTIVE" gs=0: did not work... trying in ACQUIRE mode.
05:30 pointto BS9166 TelMode="ACTIVE" gs=0; For EL =85
05:35 takepic 2.0 after collimated to < 400nm
  • image: irtc.20100629.053845.fits
05:40 pointto BS9158 TelMode="ACTIVE" gs=0; For EL=70
05:46 takepic 2.0 after collimated to < 400nm
  • image: irtc.20100629.054525.fits
05:47 pointto BS9152 TelMode="ACTIVE" gs=0; For EL=55
05:51 takepic 2.0 after collimated to < 400nm
  • image: irtc.20100629.055112.fits
05:53 pointto BS9148 TelMode="ACTIVE" gs=0; For EL=30
preset failed: no star, set into ACQUIRE mode: fixed: returned to ACTIVE mode: At low elevation, seeing is ~2.5"
06:02 takepic 2.0 after collimated to < 400nm
  • image: irtc.20100629.060225.fits
06:04 pointto BS9153 TelMode="ACTIVE" gs=0; For EL=40
06:11 takepic 2.0 after collimated to < 400nm At low elevation there is "NO collimation "available due to M2 not correcting for astigmatism. Called Runa.
  • image: irtc.20100629.061020.fits
(downtime due to extreme amounts of astigmatism)
07:09 After changing zernike values on AO system, no change was being permanently applied. Called Runa, who suggested there may be an automatic zernike override that is not allowing me to fix astigmatism manually.
07:15 Dave Thompson requests we Issue Tracked M2 problem: Issue Track # 2797

08:16 The AO team has turned off the AO automatic astigmatism.
08:16 pointto BS1969 TelMode="ACTIVE" gs=0
08:17 AO shell was rested and set flat again, and OFFSET was disabled so that the mirror would be flat and stay flat until given a zernike offset manually.
08:20 pointto BS1969 TelMode="ACQUIRE" gs=0
Seeing ~ 3"
08:22 Collimation done. Going to change elevation and hope to repeat collimation.
08:26 pointto BS9177 TelMode="ACTIVE" gs=0
08:29 Collimation done: CA = -17 IE= -78.
08:30 Seeing ~1.65". Since AO automatic astigmatism override turned off, the collimation has been working properly and consistent when changing elevation.

Preset Data

Seeing ~2.3"
Repeating Preset data on BS9166 that was taken last night: 20100628. Now at EL=45.
08:49 pointto BS9166 TelMode="ACTIVE" gs=0 PA=90
Should have put in guide star
08:50 pointto BS9166 TelMode="ACQUIRE" gs=0 PA=90 PAM-
08:52 pointto BS9166 TelMode="ACTIVE" gs=17 PA=90
09:01 Getting error in "pointto BS9166 TelMode"ACTIVE" gs=8"
  • Resmsg[0] = PresetTelescope result status: Warning
  • Resmsg[1] = Illegal command for state Failure
  • Resmsg[2] = [42] - Illegal command for state Failure
Seeing ~2.5"
09:05 takeseq 2 30
  • image: irtc.20100629.090619.fits Cube 73
09:07 offset 30 30
09:08 takeseq 2 30
  • BG image: irtc.20100629.090831.fits Cube 74

Guiding and Collimation BS9177

Seeing ~ 2.6"- 3", EL =63
09:16 pointto BS9177 Tel="ACTIVE" gs=19 PA=90
Note: The astigmatism is bad when changing from EL 45 to EL 63
IRTC> pointto BS9177 Tel="ACQUIRE" gs=0 PA=0 PAM-
Resmsg[0] = PresetTelescope result status: Warning
Resmsg[1] = Illegal command for state Failure
Resmsg[2] = [46] - Illegal command for state Failure
Changed dir to /scratch/20100629/
09:27 cl < takemanyseq.cl
  • 1st image: irtc.20100629.092733.fits Cube: 75
  • final image: irtc.20100629.095026.fits Cube: 94
Use irtc.20100629.090831.fits Cube 74 for BG

Preset Data BS9166

Returning to BS9166 for preset data at different elevation. EL =32
09:54 pointto BS9166 TelMode="ACTIVE" gs=8 PA=-90
Seeing ~3.8"
Changing guide star
09:58 pointto BS9166 TelMode="ACTIVE" gs=15 PA=90
10:02 takeseq 2 30
  • image: irtc.20100629.100256.fits Cube 95
10:04 offset 30 30
Error Messages received:
Resmsg[0] = PresetTelescope result status: Warning
Resmsg[1] = Illegal command for state Failure
Resmsg[2] = [55] - Illegal command for state Failure
Resmsg[3] = couldn't adjust exposure time to meet low counts requirement. star too dim or guide rate too high.
10:05 takeseq 2 30
  • image: irtc.20100629.100509.fits Cube 96 BG

Guiding and Collimation BS9169

Seeing ~ 3.6" -4.2": EL=42
10:09 pointto BS9169 Tel="ACTIVE" gs=14 PA=-90
10:18 cl < takemanyseq.cl
  • 1st image: irtc.20100629.101744.fits Cube 97
  • Final image: irtc.20100629.104036 Cube 116
Seeing ~ 3.2"
10:43 offset 30 30
10:44 takeseq 2 30
  • BG image: irtc.20100629.104411.fits

Twilight has started.
Going to point to BS9188 EL=70t o do a fainter guide star before closing.
10:48 pointto BS9188 Tel="ACTIVE" gs=18 PA=-90
Star not in field (also may be too faint and too bright out...)
10:52 pointto BS9188 Tel="ACQUIRE" gs=0
10:56 pointto BS9188 Tel="ACTIVE" gs=28 PA=90
Seeing ~2.2", lots of astigmatism. Marco asks to reset the active optics corrections, the image is seen to have improved by this change.
11:04 pointto BS9188 Tel="ACTIVE" gs=22 PA=-90
Collimation taking a lot time.
Seeing ~ 1.2"
11:13 cl < takemanyseq.cl
  • 1st image: irtc.20100629.111237.fits Cube 118
ABORTED sequence due to loss of guide star. May be due to twilight, may be due to collimation or faintness of star.
At EL=76 now, retrying with a brighter guide star.
11:23 pointto BS9188 Tel="ACTIVE" gs=29 PA=180
No guide star found
11:25 pointto BS9188 Tel="ACTIVE" gs=16 PA=-90
Resmsg[0] = PresetTelescope result status: Error
Resmsg[1] = Illegal command for state Failure
Resmsg[2] = [70] - Illegal command for state Failure
Resmsg[3] = Primary mirror collimation reached limit
Resmsg[4] = adjusted exposure time to 2100 ms
Resmsg[5] = guiding was not stable within 6 cycles
11:28 pointto BS9188 Tel="ACTIVE" gs=0
Resetting pointing model
11:35 pointto BS9188 Tel="ACTIVE" gs=33 PA=180
11:38 sky too bright, cannot guide. 11:41 AO shell rested.

Close

11:40 Telescope Close Humidity= 70/4%, Temp = 8 C, Wind = 5-6 m/s

-- DougMiller - 24 Jun 2010
Topic revision: r7 - 02 Jul 2010, PascaleHibon
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback